A chunk of magnesium weighing 19.4 grams and originally at 97.43 °C is dropped into an insulated cup containing 80.9 grams of water at 23.10 °C.

Assuming that all of the heat is transferred to the water, the final temperature of the water is °C.
Answer #1

Answer

Heat lost by magnesium = Heat gained by water

mass of Magnesium *specific feat of magnesium* temperature change = mass of water**specific feat of water* temperature change

19.4 grams*1.02 J/gC* (97.43- Tf)= 80.9 grams*4.184 J/gC*(Tf-23.10)

19.788*(97.43- Tf)  = 338.48 *(Tf-23.10)

1927.94-19.788Tf = 338.48Tf- 7818.88

1927.94+7818.88 = 338.48Tf+19.788Tf

9746.828 =358.268Tf

Tf =9746.828/358.268

Tf = 27.2 degree C